question C: im done eoth a and c. im not sure about c
421 On August 3, Cinco Construction purchased special-purpose equipment at a cost of $1,000,000. The useful life of the equipment was estimated to be eight years, with an estimated residual value of $50,000. S a. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the straight-line depreciation method (half-year convention). b. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the 200 percent declining-balance method (half-year convention) with a switch to straight-line when it will maximize depreciation expense. c. Which of these two depreciation methods (straight-line or double-declining-balance) results in the highest net income for financial reporting purposes during the first two years of the equipment's use? Explain. On January 2, 2018, Jatson Corporation acquired five years. The cost of the equipment a. Prepare a complete depreciation table under the three depreciation methods listed as follows. a new machine with an estimated useful life of was $40,000 with an estimated residual value of $5,000. S
421 On August 3, Cinco Construction purchased special-purpose equipment at a cost of $1,000,000. The useful life of the equipment was estimated to be eight years, with an estimated residual value of $50,000 Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the straight-line depreciation method (half-year convention). b. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the 200 percent declining-balance method (half-year convention) with a switch to straight-line when it will maximize depreciation expense. Which of these two depreciation methods (straight-line in the highest net income for financial reporting purposes during the first two years of the equipment's use? Explain. a. double-declining-balance) results or c. Jatson Corporation acquired a new machine with an estimated useful life of
Answer #1

Part A

Annual depreciation = (cost – salvage value)/estimated life = (1000000-50000)/8 = 118750

For half year it will be = 118750/2 = 59375

year

Straight-Line (Half-Year Convention)

1

59375

2

118750

3

118750

4

118750

5

118750

6

118750

7

118750

8

118750

9

59375

Part B

Depreciation rate = 1/8 * 2 = 25%

Depreciation = book value of each * depreciation rate

Book value = previous year’s carrying value – current year’s depreciation

year

Straight-Line (Half-Year Convention)

1

125000

2

218750

3

164063

4

123047

5

92285

6

69214

7

51910

8

38933

9

66798

year

Beginning value

Depreciation

Carrying value

0

1000000

1

1000000

125000

875000

2

875000

218750

656250

3

656250

164063

492188

4

492188

123047

369141

5

369141

92285

276855

6

276855

69214

207642

7

207642

51910

155731

8

155731

38933

116798

9

116798

66798

50000

Part C

Straight line method will results in the highest net income as the depreciation amount in the straight line method is low in the year 1 and year 2.
Answer #2

B. This is wrong in the last 3 Depreciations. You take the last book value of 207641 and use that calculation to go to straight line depreciation since it will be below the residual value.

Switch to straight-line: ($207,641 – $50,000) ÷ 3 years (that are left in the schedule) = $52,547

answered by: Joe Shmoe